# Leveraging Multimodal LLM for Inspirational User Interface Search

### v2

@@ 4.2.1. Search Query and Retrieval Method Design
We designed a query system that enables users to input their desired semantics and retrieve the most relevant UI designs from the database. The search query input consists of user-specified semantics, where users can select the semantics they want to search for and provide a textual description for each semantic, enabling a flexible combination of functional requirements and aesthetic intent (DP1). To measure the similarity between the user query and the UI designs in the database, we employed the OpenAI text-embedding-3-large model for text embedding. As shown in Figure 3, we calculated the cosine similarity between the query and semantics in the database. Guided by the principle of flexible queries for adjusting priorities (DP2), in cases where multiple semantics are involved, users can adjust the weight of each semantic to reflect their importance, and a weighted sum is applied to calculate the overall similarity score. To construct the search database, we extracted text embeddings for all UI semantics in each screenshot. For semantics in the key-value format in the YAML output, we testified the semantics by including the key along with the value description. This approach ensures that the similarity search not only considers the textual description in the value but also reflects the specific semantic represented by the key. Furthermore, aligned with our principle of supporting functionality aspects of UI (DP1), we developed special queries to address this requirement. While the regular queries focus on finding UI designs similar to the user-specified semantics, these special queries are designed to retrieve the next or previous screens in the user flow. This functionality is made possible by the LLM’s ability to infer the next and previous screens. The special queries are implemented by calculating the similarity between the screen roles in the database and the inferred next or previous screen semantics from the query.
@@ 4.2.2. Interactive Search System
The search panel enables users to define their desired semantics, providing a customizable blend of functional needs and visual preferences (DP1). Users can freely add or remove semantic attributes, input natural language descriptions for specific design requirements, and utilize dropdowns for predefined categorical semantics. To reflect designers’ needs to adjust priorities (DP2), the system allows users to assign weights to each semantic, visually represented by a dynamic bar chart above the search button. This functionality helps designers fine-tune their queries and prioritize key aspects of UI design based on their needs. Once the query is submitted, the backend processes the selected semantics and weights to retrieve the most relevant UI designs from the database. The system uses pre-computed text embeddings of the UI semantics for efficient similarity calculation and ranking.
The results panel presents retrieved UI designs in a structured format, with overall matches considering all provided semantics at the top. Below, individual semantic search results are shown, ranked by similarity scores for each facet. Supporting our principle of exploratory and adaptive search (DP3), clicking on a result image opens a screen detail panel that displays the extracted semantics for the screen. This detailed view, as shown in the bottom part of  Figure 4, allows designers to examine the semantic attributes in-depth, assessing the design’s suitability for their project. To facilitate the iterative nature of design exploration, the import button in the detailed view enables users to incorporate the semantics into subsequent queries, supporting an iterative and exploratory design process. We also integrated this feature into the search panel to support a query-by-example approach for exploratory searches, enabling users to upload their own UI images. Upon image upload, the system employs the MLLM to extract semantics, which designers can review and import into the search panel to create structured semantic queries.
@@ 5. Computational Evaluation
To assess the effectiveness of our proposed method, we conducted a two-step evaluation process. In the first step, we quantitatively measured UI semantic understanding capability that could be objectively measured using existing UI datasets. In the second step, detailed in Section 6, we conducted human evaluations with UI designers, focusing on two aspects: (1) assessing the quality of extracted UI semantics and (2) performing a comparative study where designers conducted searches for design inspiration using our retrieval system S&UI and an existing tool. This dual evaluation approach enabled us to understand how the extracted semantics influence the overall search experience.
@@
Figure 5. Comparison of Correct Predictions: The top chart compares GPT-4o and GUIClip across 20 screen categories, while the bottom chart contrasts GPT-4o and CLIP over 31 app categories. Top-1 predictions are highlighted, with top-3 shown semi-transparently. GPT-4o outperforms baseline methods in both cases.
@@
5.1. App and Screen Category Classification
As shown in Table 4, GPT-4o significantly outperformed the baseline models in both screen category and app category classification, achieving a top-1 accuracy of 59.21% and 58.33%, respectively, compared to the baselines, which achieved 36.24% and 39.95%. Moreover, GPT-4o’s top-3 accuracy reached 78.83% for screen categories and 75.26% for app categories, outperforming the baseline performances of 58.32% and 60.57%. These results demonstrate that MLLMs can achieve even stronger UI semantic understanding than vision foundation models like CLIP, which have already established significant improvements over traditional UI-specific approaches. To understand the performance result more deeply, we conducted a detailed category-wise experiment comparing with the most substantial baseline for each task, as shown in Figure 5. The results reveal that GPT-4o outperforms in most categories. Interestingly, GUIClip showed higher accuracy in some categories like TUTORIAL and SEARCH; we assume that this is likely due to labeling ambiguities in the Enrico dataset. TUTORIAL is explained as “Onboarding screen” can be confused with BARE and LOGIN screens, likewise the “Search engine functionality” label of SEARCH may not accurately represent basic search interface screens. Despite these exceptions, the results demonstrate that MLLMs can achieve more substantial UI screen and application understanding than baselines.
@@ 5.2. UI Element Prediction
We used the CLAY dataset (Li et al., 2022), which contains annotated UI elements for a subset of the Rico screenshots. As we focus on inspirational search rather than precise element bounding box prediction, we evaluated the model’s ability to identify the presence of different UI element types within a screen. We provided GPT-4o with the CLAY UI element labels and descriptions, instructing the model to extract only the specified UI element semantics in a zero-shot manner. We excluded the ROOT, BACKGROUND, and CONTAINER labels from the evaluation as they do not represent typical UI elements, and also excluded PICTOGRAM as the label is confusing with icon and illustration. The evaluation was conducted on the 1,318 screenshots present in both the Enrico and CLAY datasets. GPT-4o achieved a weighted average precision of 0.661, recall of 0.703, and F1-score of 0.681 across all UI element types. All metrics are weighted by the number of each label in the dataset.
@@ 5.3. Screen Role Understanding
However, the potential of MLLMs in extracting screen descriptions using images has not been investigated. Thus, we compare the screen role (description) extracted by GPT-4o with the Screen2Words dataset to evaluate the richness of the generated explanations in a zero-shot manner. To assess the quality of the screen descriptions generated by GPT-4o, we calculated the syntactic dependency complexity and Part-of-Speech (POS) diversity scores using spacy (Honnibal et al., 2020). Syntactic dependency complexity measures the intricacy of the syntactic structure, while POS diversity evaluates the variety of word types used in the descriptions. As the Screen2Words dataset contains five descriptions per screen, we compared the average scores and the highest-scoring items from the Screen2Words with the screen role extracted by GPT-4o, as shown in Figure 6.
The results demonstrate that GPT-4o generates more complex and diverse text compared to the average and best-performing items in the Screen2Words dataset. The higher dependency complexity and POS diversity scores indicate that GPT-4o captures a broader range of syntactic structures and word types, leading to more informative screen roles. Our findings highlight the powerful capability of MLLMs to extract screen roles or descriptions. By leveraging visual information from UI screens, MLLM can generate richer and more comprehensive descriptions than the Screen2Words dataset.
@@ 6. Human Evaluation
To assess the quality and usefulness of the semantics extracted by the MLLM from human perspectives (RQ2) and evaluate the advantages of our semantic-based retrieval system (RQ3), we conducted a two-part human evaluation with UI designers. To address RQ2, Section 6.2 involves UI designers assessing the quality of extracted semantics across 14 categories. This offers insight into the quality and usefulness of the MLLM-extracted semantics. Section 6.3 tackles RQ3 by employing both our semantic-based search tool and a baseline for realistic inspiration-finding tasks. Analyzing the ratings and feedback from both systems reveals the advantages and user experience benefits of our approach.
@@ 6.1. Participants
We recruited 10 participants (4 male, 6 female, denoted E1 - E10) from designer communities and through word-of-mouth, including students, novice designers, and those with professional design experiences, as detailed in Table 5. The number of participants and their diverse expertise are considered to align with formative study and practices in computational UI research, where similar studies have included around 10 participants and varying levels of expertise (Wu et al., 2024a, b). The participants took part in both the 30-minute quality assessment and 1-hour comparative study of retrieval method evaluation, allowing for integrated insights on the semantics and retrieval system. Participants were compensated at a rate equivalent to USD 22 for the 1.5-hour study session. The study received approval from the Institutional Review Board (IRB).
@@ 6.2.1. Quantitative Assessment of Semantics
Notably, the serendipity scores exhibit a different distribution than relevance and comprehensiveness. Categories like app description and screen description, which focus on well-established information, received lower serendipity scores, indicating that GPT-4o provided highly accurate yet expected results in these areas. In contrast, semantics related to similar app, target user, and screen flow (next screen and previous screen), which surface more unique and less commonly encountered information, scored higher on serendipity (4.6$\pm$1.57, 4.5$\pm$1.17, 4.6$\pm$2.06, and 4.3$\pm$1.88 respectively). This suggests that MLLM excelled at uncovering unexpected yet valuable information in cases where designers might not have initially considered such details, enhancing its potential for inspiring novel design ideas. On the other hand, the color scheme semantic received the lowest serendipity score (2.6$\pm$1.50), suggesting that the extracted color schemes align more closely with designers’ expectations and provide fewer unexpected ideas.
Comparing across the different levels, we observe that the app and screen level semantics generally received higher relevance and comprehensiveness scores compared to the composition and visual design levels. This suggests that MLLM is particularly effective at extracting high-level, functional information about apps and screens. However, the serendipity scores are more evenly distributed across levels, with the composition level semantics (e.g., action items) showing promise for surfacing novel design ideas.
@@ 6.2.2. Qualitative Interview Insights on Extracted Semantics
To gain a deeper understanding of designers’ perceptions and experiences with the extracted semantics, we conducted semi-structured interviews following the quantitative evaluation. We asked participants for their impressions of each semantic across levels, probing aspects such as accuracy, usefulness, limitations, and potential improvements. We analyzed the interview responses using thematic analysis, identifying recurring themes and patterns across participants’ feedback for each semantic level. The interviews provided rich insights into each level’s strengths, limitations, and potential use cases of the extracted semantics.
App Level Semantics
Participants generally found the app-level semantics (i.e. app category, description, and target user) accurate and informative. They appreciated the level of detail provided in the app descriptions and target user information, which often went beyond their initial expectations (E1-3, 5, 10). For example, E10 highlighted a case where the target user description for a banking app included specific details about users’ concerns for privacy and security. However, some participants (E1, 4, 8) mentioned that similar app suggestions were sometimes repetitive or included well-known apps, making it difficult to assess their serendipity. Participants also noted that while the app category and description information was accurate, it might not always be necessary or useful in their design process (E5, 8). Several participants were impressed by how the model could extract meaningful application-level information from latent details in single UI screenshots, such as background images or small text elements (E2, 7, 10).
@@ 6.3.2. Qualitative Interview Insights on S&UI
Several participants highlighted the value of the iterative search process enabled by S&UI. They appreciated how the system allowed them to refine their queries based on previous result screens and their semantics descriptions, helping them clarify their search strategies about design intentions and discover new ideas along the way. E8 described it as “narrowing down the search step by step,” while E3 noted, ”The ability to iteratively explore and refine my search was really helpful in finding the right inspiration.” This iterative approach was less evident in the baseline system, which often required users to start new searches from scratch when refining their queries.
@@
We found interesting patterns in how designers with varying experience levels utilized S&UI. Junior designers (3̃ years experience) approached the tool to understand foundational design elements (E2, 5, 6), as E6 noted: “I search for screens with a similar function because it helps me figure out the layout and elements I need”. Senior designers (5+ years experience) leveraged more sophisticated search strategies (E3, 7, 9, 10), with E7 stating: “I combined semantic features together to find exactly what I needed”. This difference in approach was particularly evident when compared to traditional tools, as highlighted by E3, an experienced designer: “Current tools only let you search with keywords, but they don’t give you the deeper context, like the app’s audience or the screen’s role”. These findings indicate that S&UI supports both learning and professional workflows through its semantic-based approach. This adaptability across experience levels suggests that semantic-based UI search tools can provide meaningful support while accommodating different needs and expertise.
@@ 7. Discussion and Future Work
Our study leverages MLLMs to extract UI semantics from mobile UI images, enhancing design inspiration search and surpassing existing methods. This section explores MLLMs’ UI semantic understanding, how we enhanced the design inspiration search process, implications for UI design search tools, and the release of the S&UI dataset, while discussing limitations and future directions.
@@ 7.1. UI Semantic Understanding with MLLM
We found that MLLMs can effectively extract meaningful semantics from UI designs. According to computational evaluation, MLLM shows strength in multiple UI semantic understanding, such as app categories and screen roles, significantly outperforming existing approaches. Likewise, designers appreciated the model’s ability to extract those semantics accurately. They also appreciated our unique semantics, which had previously been unavailable, like target users, similar apps, screen flows, and mood. These semantics earned relatively higher serendipity ratings. Meanwhile, we observed variations in performance across different semantic levels. In the human evaluation, app and screen level semantics achieved high relevance and comprehensiveness ratings from designers, while composition and visual design level semantics showed few limitations due to their complexity and subjectivity. The model’s extraction capabilities correlate with the concreteness of semantics. The MLLM performs best when extracting categorical (e.g., app and screen categories) and descriptive semantics (e.g., screen role, app description), but struggles with complex, multiple attributes like layouts, UI elements, and visual design aspects.

### v3

@@ 4.2.1. Search Query and Retrieval Method Design
We designed a query system that enables users to input their desired semantics and retrieve the most relevant UI designs from the database. The search query input consists of user-specified semantics, where users can select the semantics they want to search for and provide a textual description for each semantic, enabling a flexible combination of functional requirements and aesthetic intent (DP1). To measure the similarity between the user query and the UI designs in the database, we employed the OpenAI text-embedding-3-large model for text embedding. As shown in Figure 3, we calculated the cosine similarity between the query and semantics in the database. Guided by the principle of flexible queries for adjusting priorities (DP2), in cases where multiple semantics are involved, users can adjust the weight of each semantic to reflect their importance, and a weighted sum is applied to calculate the overall similarity score. To construct the search database, we extracted text embeddings for all UI semantics in each screenshot. For semantics in the key-value format in the YAML output, we represented the semantics by including the key along with the value description. This approach ensures that the similarity search not only considers the textual description in the value but also reflects the specific semantic represented by the key. Furthermore, aligned with our principle of supporting functionality aspects of UI (DP1), we developed special queries to address this requirement. While the regular queries focus on finding UI designs similar to the user-specified semantics, these special queries are designed to retrieve the next or previous screens in the user flow. This functionality is made possible by the LLM’s ability to infer the next and previous screens. The special queries are implemented by calculating the similarity between the screen roles in the database and the inferred next or previous screen semantics from the query.
@@ 4.2.2. Interactive Search System
The search panel enables users to define their desired semantics, providing a customizable blend of functional needs and visual preferences (DP1). Users can freely add or remove semantic attributes, input natural language descriptions for specific design requirements, and utilize dropdowns for predefined categorical semantics. To adjust search priorities (DP2), the system allows users to assign weights to each semantic, visually represented by a dynamic bar chart above the search button. This functionality helps designers fine-tune their queries and prioritize specific UI semantics. Once the query is submitted, the backend processes the selected semantics and weights to retrieve the most relevant UI designs from the database. The system uses pre-computed text embeddings of the UI semantics for efficient similarity calculation and ranking.
The results panel presents retrieved UI designs in a structured format, with overall matches considering all provided semantics at the top. Below, individual semantic search results are shown, ranked by similarity scores for each facet. To support exploratory and adaptive search (DP3), clicking on a result image opens a screen detail panel that displays the extracted semantics for the screen. This detailed view, as shown in the bottom part of  Figure 4, allows designers to examine the semantic attributes in-depth, assessing the design’s suitability for their project. To facilitate the iterative nature of design exploration, the import button in the detailed view enables users to incorporate the semantics into subsequent queries, supporting an iterative and exploratory design process. We also integrated this feature into the search panel to support a query-by-example approach for exploratory searches, enabling users to upload their own UI images. Upon image upload, the system employs the MLLM to extract semantics, which designers can review and import into the search panel to create structured semantic queries.
@@ 5. Computational Evaluation
To assess the effectiveness of our proposed method, we conducted a two-step evaluation process. In the first step, we quantitatively measured UI semantic understanding capability that could be objectively measured using existing UI datasets. In the second step, detailed in Section 6, we conducted human evaluations with UI designers, focusing on two aspects: (1) assessing the quality of extracted UI semantics and (2) conducting a comparative study where designers conducted searches for design inspiration using our retrieval system S&UI and an existing tool. This dual evaluation approach allowed us to understand how the extracted semantics influence the overall search experience.
@@ 5.1. App and Screen Category Classification
Figure 5. Comparison of Correct Screen and App Category Predictions: The top chart compares GPT-4o and GUIClip across 20 screen categories, while the bottom chart contrasts GPT-4o and CLIP over 31 app categories. Top-1 predictions are highlighted, with top-3 shown semi-transparently. GPT-4o outperforms baseline methods in both cases.
@@
As shown in Table 4, GPT-4o significantly outperformed the baseline models in both screen category and app category classification, achieving a top-1 accuracy of 59.21% and 58.33%, respectively, compared to the baselines, which achieved 36.24% and 39.95%. Moreover, GPT-4o’s top-3 accuracy reached 78.83% for screen categories and 75.26% for app categories, outperforming the baseline performances of 58.32% and 60.57%, respectively. These results demonstrate that MLLMs can achieve even stronger UI semantic understanding than vision foundation models like CLIP, which have already established significant improvements over traditional UI-specific approaches. To understand the performance result more deeply, we conducted a detailed category-wise experiment comparing with the most substantial baseline for each task, as shown in Figure 5. The results reveal that GPT-4o outperforms in most categories. Interestingly, GUIClip showed higher accuracy in some categories like TUTORIAL and SEARCH; we assume that this is likely due to labeling ambiguities in the Enrico dataset. TUTORIAL is explained as “Onboarding screen” can be confused with BARE and LOGIN screens, likewise the “Search engine functionality” label of SEARCH may not accurately represent basic search interface screens. Despite these exceptions, the results demonstrate that MLLMs can achieve more substantial UI screen and application understanding than baselines.
@@ 5.2. UI Element Prediction
We used the CLAY dataset (Li et al., 2022), which contains annotated UI elements for a subset of the Rico screenshots. As we focus on inspirational search rather than precise element bounding box prediction, we evaluated the model’s ability to identify the presence of different UI element types within a screen. We provided GPT-4o with the CLAY UI element labels and descriptions, instructing the model to extract only the specified UI element semantics in a zero-shot manner. We excluded the ROOT, BACKGROUND, and CONTAINER labels from the evaluation as they do not represent typical UI elements, and also excluded PICTOGRAM as the label is confusing with icon and illustration. The evaluation was conducted on the 1,318 screenshots present in both the Enrico and CLAY datasets. GPT-4o achieved a weighted average precision of 0.661, recall of 0.703, and $F_{1}$-score of 0.681 across all UI element types. All metrics are weighted by the number of each label in the dataset.
@@ 5.3. Screen Role Understanding
However, the potential of MLLMs in extracting screen descriptions using UI screen images has not been investigated. Thus, we compare the screen role (description) extracted by GPT-4o with the Screen2Words dataset to evaluate the richness of the generated explanations in a zero-shot manner. To assess the quality of the screen descriptions generated by GPT-4o, we calculated the syntactic dependency complexity and Part-of-Speech (POS) diversity scores using spacy (Honnibal et al., 2020). Syntactic dependency complexity measures the intricacy of the syntactic structure, while POS diversity evaluates the variety of lexical types used in the descriptions. As the Screen2Words dataset contains five descriptions per screen, we compared the average scores and the highest-scoring items from the Screen2Words with the screen role extracted by GPT-4o.
As shown in Figure 6, the results indicate that GPT-4o generates more complex and diverse descriptive text compared to both the average and highest-performing items in the Screen2Words dataset. The higher scores for syntactic dependency complexity and POS diversity indicate that GPT-4o effectively captures a wider range of syntactic structures and richer lexical repertoires, which leads to more informative and exhaustive screen descriptions. Our findings suggest that MLLM can generate more detailed and comprehensive screen descriptions than traditional methods by utilizing visual information from UI screen images.
@@ 6. Human Evaluation
To assess the quality and usefulness of the semantics extracted by the MLLM from human perspectives (RQ2) and evaluate the advantages of our semantic-based retrieval system (RQ3), we conducted a two-part human evaluation with UI designers. To address RQ2, Section 6.2 involves UI designers assessing the quality of each extracted semantics. This offers insight into the quality and usefulness of the MLLM-extracted semantics. Section 6.3 tackles RQ3 by employing both our semantic-based search tool and a baseline for realistic inspiration-finding tasks. Analyzing the ratings and feedback from both systems reveals the advantages and user experience benefits of our approach.
@@ 6.1. Participants
We recruited 10 participants (4 male, 6 female, denoted E1 - E10) from designer communities and through word-of-mouth, including students, novice designers, and those with professional design experiences, as detailed in Table 5. The number of participants and their diverse expertise are considered to align with formative study and practices in computational UI research, where similar studies have included around 10 participants and varying levels of expertise (Wu et al., 2024a, b). The participants took part in both the 30-minute quality assessment and 1-hour comparative study of retrieval method evaluation, allowing for integrated insights into the semantics and retrieval system. Participants were compensated at a rate equivalent to USD 22 for the 1.5-hour study session. The study received approval from the Institutional Review Board (IRB).
@@ 6.2.1. Quantitative Assessment of Semantics
Notably, the serendipity scores exhibit a different distribution than relevance and comprehensiveness. Categories like app description and screen description, which focus on well-established information, received lower serendipity scores, indicating that GPT-4o provided highly accurate yet expected results in these areas. In contrast, semantics related to similar app, target user, and screen flow (next screen and previous screen), which surface more unique and less commonly encountered information, scored higher on serendipity (4.6$\pm$1.58, 4.5$\pm$1.18, 4.6$\pm$2.07, and 4.3$\pm$1.89 respectively). This suggests that MLLM excelled at uncovering unexpected yet valuable information in cases where designers might not have initially considered such details, enhancing its potential for inspiring novel design ideas. On the other hand, the color scheme semantic received the lowest serendipity score (2.6$\pm$1.50), suggesting that the extracted color schemes align more closely with designers’ expectations and provide fewer unexpected ideas.
Comparing across the different levels, we observed that the app and screen level semantics generally received higher relevance and comprehensiveness scores compared to the composition and visual design levels. This suggests that MLLM is particularly effective at extracting high-level, functional information about apps and screens. However, the serendipity scores are more evenly distributed across levels, with the composition level semantics (e.g., action items) showing promise for surfacing novel design ideas.
@@ 6.2.2. Qualitative Interview Insights on Extracted Semantics
To gain a deeper understanding of designers’ perceptions of the extracted semantics, we conducted semi-structured interviews following the quantitative evaluation. We asked participants for their impressions of each semantic across levels, probing aspects such as accuracy, usefulness, limitations, and potential improvements. We analyzed the interview responses using thematic analysis, identifying recurring themes and patterns across participants’ feedback for each semantic level. The interviews provided rich insights into each level’s strengths, limitations, and potential use cases of the extracted semantics.
Application Level Semantics
Participants generally found the app-level semantics (i.e., app category, description, and target user) accurate and informative. They appreciated the level of detail provided in the app descriptions and target user information, which often went beyond their initial expectations (E1-3, 5, 10). For example, E10 highlighted a case where the target user description for a banking app included specific details about users’ concerns for privacy and security. However, some participants (E1, 4, 8) mentioned that similar app suggestions were sometimes repetitive or included well-known apps, making it difficult to assess their serendipity. Participants also noted that while the app category and description information was accurate, it might not always be necessary or useful in their design process (E5, 8). Several participants were impressed by how the model could extract meaningful application-level information from latent details in single UI screenshots, such as background images or small text elements (E2, 7, 10).
@@ 6.3.2. Qualitative Interview Insights on S&UI
Several participants highlighted the value of the iterative search process enabled by S&UI. They appreciated how the system allowed them to refine their queries based on previous result screens and their semantics descriptions, helping them clarify their search strategies about design intentions and discover new ideas along the way. E8 described it as “narrowing down the search step by step,” while E3 noted, “The ability to iteratively explore and refine my search was really helpful in finding the right inspiration.” This iterative approach was less evident in the baseline system, which often required users to start new searches from scratch when refining their queries.
@@
We found interesting patterns in how designers with varying experience levels utilized S&UI. Junior designers (3- years of experience) approached the tool to understand foundational design elements (E2, 5, 6), as E6 noted: “I search for screens with a similar function because it helps me figure out the layout and elements I need”. Senior designers (5+ years of experience) leveraged more sophisticated search strategies (E3, 7, 9, 10), with E7 stating: “I combined semantic features together to find exactly what I needed”. This difference in approach was particularly evident when compared to traditional tools, as highlighted by E3, an experienced designer: “Current tools only let you search with keywords, but they don’t give you the deeper context, like the app’s audience or the screen’s role”. These findings indicate that S&UI supports both learning and professional workflows through its semantic-based approach. This adaptability suggests that semantic-based UI search tools can provide meaningful support while accommodating different needs and expertise.
@@ 7. Discussion and Future Work
Our study leverages MLLM to extract UI semantics from mobile UI images, enhancing design inspiration search and surpassing existing methods. This section explores MLLM’s UI semantic understanding, how we enhanced the design inspiration search process, implications for UI design search tools, and the release of the S&UI dataset, while discussing limitations and future directions.
@@ 7.1. UI Semantic Understanding with MLLM
We found that MLLMs can effectively extract meaningful semantics from UI designs. According to computational evaluation, MLLM shows strength in multiple UI semantic understandings, such as app categories and screen roles, significantly outperforming existing approaches. Likewise, designers appreciated the model’s ability to extract those semantics accurately. They also appreciated our unique semantics, which had previously been unavailable, like target users, similar apps, screen flows, and mood. These semantics earned relatively higher serendipity ratings. Meanwhile, we observed variations in performance across different semantic levels. In the human evaluation, app and screen level semantics achieved high relevance and comprehensiveness ratings from designers, while composition and visual design level semantics showed few limitations due to their complexity and subjectivity. The model’s extraction capabilities correlate with the concreteness of semantics. The MLLM performs best when extracting categorical (e.g., app and screen categories) and descriptive semantics (e.g., screen role, app description), but struggles with complex, multiple attributes like layouts, UI elements, and visual design aspects.